# TDS-CLIP: Temporal Difference Side Network for Image-to-Video Transfer Learning

### v1

@@ Abstract
Recently, large-scale pre-trained vision-language models (e.g., CLIP), have garnered significant attention thanks to their powerful representative capabilities. This inspires researchers in transferring the knowledge from these large pre-trained models to other task-specific models, e.g., Video Action Recognition (VAR) models, via particularly leveraging side networks to enhance the efficiency of parameter-efficient fine-tuning (PEFT). However, current transferring approaches in VAR tend to directly transfer the frozen knowledge from large pre-trained models to action recognition networks with minimal cost, instead of exploiting the temporal modeling capabilities of the action recognition models themselves. Therefore, in this paper, we propose a memory-efficient Temporal Difference Side Network (TDS-CLIP) to balance knowledge transferring and temporal modeling, avoiding backpropagation in frozen parameter models. Specifically, we introduce a Temporal Difference Adapter (TD-Adapter), which can effectively capture local temporal differences in motion features to strengthen the model’s global temporal modeling capabilities. Furthermore, we designed a Side Motion Enhancement Adapter (SME-Adapter) to guide the proposed side network in efficiently learning the rich motion information in videos, thereby improving the side network’s ability to capture and learn motion information. Extensive experiments are conducted on three benchmark datasets, including Something-Something V1&V2, and Kinetics-400. Experimental results demonstrate that our approach achieves competitive performance. Code will be available at https://github.com/BBYL9413/TDS-CLIP.
@@
In recent years, Large Language Models (LLMs) have achieved significant success (Mann et al. 2020; Touvron et al. 2023; Ryoo et al. 2021) in the fields such as text-video retrieval (Wu et al. 2023a), text-to-image generation (Qu et al. 2023), etc., gradually extending their influence to the domain of computer vision. Video Action Recognition (VAR) tasks, a challenging research area within computer vision, have seen mainstream large language models represented by the CLIP paradigm (Radford et al. 2021), which includes a Vision Transformer (ViT) vision encoder (Dosovitskiy et al. 2020) for image processing and a text encoder. CLIP establishes a strong association between images and text and is applied to various video tasks. This raises a thought-provoking question: how can we effectively utilize the pre-training of such large models for VAR tasks?
@@
Figure 1: Comparison of per-video GFLOPs and accuracy on SSv2 (Goyal et al. 2017). The size of the bubbles represents the memory usage during training. EVL: Efficient Video Learning (Lin et al. 2022). ST-Adapter: Spatio-Temporal Adapter (Pan et al. 2022).
@@
Recently, the emergence of Parameter-Efficient Fine-Tuning (PEFT) methods (Houlsby et al. 2019; Lester, Al-Rfou, and Constant 2021) has effectively alleviated this issue. PEFT adapts specific tasks by adjusting the learnability of certain parameters or introducing various adapters, avoiding extensive parameter adjustments across the entire model. With the advent of PEFT, researchers are exploring the possibility of freezing the original CLIP parameters and incorporating various adapters (Wang et al. 2024) or prompts (Wasim et al. 2023). These methods achieve impressive results in video understanding and are widely used in tasks such as video action recognition. As shown in Fig. 2(a), a popular design for efficient transfer learning is to insert tunable structures between pre-trained and parameter-frozen CLIP visual transformer blocks. Although introducing adapters under frozen parameters facilitates temporal action modeling, backpropagation still passes through the frozen layers, consuming memory. To enhance memory efficiency, LST (Sung, Cho, and Bansal 2022) introduces a side network connected to the frozen pre-trained model for backpropagation, as shown in Fig. 2(b). Continuing with this idea, Side4Video (Yao et al. 2023) further explores the capability of side networks in temporal modeling within the computer vision domain, achieving parameter-efficient video understanding tasks. These methods treat the spatial and temporal branches equally. However, for VAR tasks, it is often necessary to emphasize the effective processing of the temporal branch. The potential of using side networks for temporal modeling in video understanding tasks remains to be fully explored.
@@
Figure 2: Comparison with commonly used and effective fine-tuning methods for video recognition. (a): ST-Adapter: Spatio-Temporal Adapter (Pan et al. 2022). (b): Ladder Side-tuning (Sung, Cho, and Bansal 2022).
@@
In this work, we propose a method based on CLIP fine-tuning that balances memory efficiency and temporal modeling performance, called TDS-CLIP. First, we shift the adapter positions used for the CLIP vision branch to the side network to achieve memory-efficient pre-training fine-tuning. To better model the temporal information of videos, we design a new Temporal Difference Adapter (TD-Adapter) that effectively captures local temporal differences to enhance the model’s global temporal modeling capability. Second, we integrate the TD-Adapter with a multi-head self-attention module and feed-forward network to form a learnable side network module, used for parameter training and backpropagation. Finally, to improve the side network’s capability of capturing video motion information, we introduce a Side Motion Enhancement Adapter (SME-Adapter) that acts on the input of the side network to guide it in effectively learning additional motion information relevant to action labels.
@@
We evaluate the performance of our method on several large-scale temporal action recognition datasets, including Something-Something V1(Goyal et al. 2017), Something-Something V2(Goyal et al. 2017), and Kinetics-400(Kay et al. 2017). Experimental results demonstrate that the proposed method effectively accomplishes video action understanding tasks.
@@
  • We propose a novel memory-efficient Temporal Difference Side Network (TDS-CLIP) for video action recognition that balances knowledge transfer and temporal modeling without requiring backpropagation in models with frozen parameters
@@
VLM for Video Recognition. ViT (Dosovitskiy et al. 2020) models are gradually replacing convolutional neural networks in the mainstream due to their powerful learning capabilities and flexibility. EVA-CLIP (Sun et al. 2023) proposes a 64-layer ViT-E with larger scale parameters, and the effective adaptation of such a huge image model to the video domain is a task well worth exploring. Meanwhile, Transferring pre-trained models of language-images from the ViT architecture to video attracts widespread attention (Cheng et al. 2021). Previous action recognition methods rely on pre-trained models from ImageNet (Simonyan and Zisserman 2014). Currently, powerful text-image pre-trained models based on CLIP are gradually emerging and are increasingly used in VAR tasks (Wang et al. 2023). EVL (Lin et al. 2022) uses parallel Transformer decoders to extract spatial features from the frozen CLIP model for VAR tasks. ST-Adapter (Pan et al. 2022) introduces a spatiotemporal convolution adapter to endow the CLIP model with the capability of understanding videos. AIM (Yang et al. 2023) introduces spatial adaptation, temporal adaptation, and joint adaptation, gradually equipping image models with spatiotemporal reasoning abilities, further enriching the types of adapters to suit different video action recognition tasks. DUALPATH (Park et al. 2023) achieves dynamic modeling using spatial and temporal adaptation paths, merging successive frames into a grid-like frameset to accurately infer relationships between markers. M2-CLIP (Wang et al. 2024) introduces a TED-Adapter, enhancing the model’s generalization ability while maintaining video action recognition performance, and successfully uses it for multimodal VAR tasks. These methods effectively transfer the powerful pre-trained CLIP visual encoder to VAR tasks but do not address the issue of memory consumption during backpropagation.
@@
Side Network. The side-tuning network (Zhang et al. 2020) is first proposed to address overfitting and the problem of forgetting in incremental learning, and it applies to various downstream tasks. In vision tasks, LST (Sung, Cho, and Bansal 2022) tackles this problem by implementing a side branch network attached to the CLIP visual model. This scheme is effective in avoiding the passing of gradients in the visual encoder of frozen parameters thus reducing memory usage. Inspired by the LST, DiST (Qing et al. 2023) uses a dual encoder structure, employing the pre-trained CLIP base model as the spatial encoder and introducing a lightweight additional network as the temporal encoder, finally integrating spatiotemporal information through an ensemble branch to achieve memory-efficient video understanding. Side4Video (Yao et al. 2023) further explores the effective forms of memory-efficient side networks in video action recognition tasks and proposes a spatiotemporal side network for video understanding tasks. Although these methods demonstrate that using side networks can achieve memory-efficient image-to-video transfer learning, they do not deeply explore the temporal modeling capabilities of side network models, which we further enhance. In this work, we introduce corresponding temporal modeling adapters at both the input and the path of the side network composed of ViT encoders, transforming it into a memory-efficient and powerful temporal modeling framework for VAR tasks.
Method
@@
As shown in Fig. 3, the overall architecture of TDS-CLIP framework consists of two key components: the visual spatial encoder and the visual temporal encoder. The visual spatial encoder is composed of a frozen CLIP pre-trained ViT, which provides strong spatial semantic features for the input video frames. The visual temporal encoder is a side network structure that accepts the output features from the visual spatial encoder. It takes the dense video frame features provided by the SME-Adapter as input and transforms the spatial features from the frozen CLIP branch into an effective form suitable for temporal modeling in video understanding.
@@
Visual Spatial Encoder
The spatial encoder consists of the ViT with frozen parameters from the CLIP visual branch, which can independently capture the spatial semantic features of sparse input video frames, thereby leveraging its powerful pre-training to enhance training performance. Following the ViT format, the input frames are first divided into $N=\frac{HW}{P^{2}}$ patches, each with a shape of $P^{2}$, and then projected through a fully connected layer. Then, for the input full-length video frame sequence, each patch feature vector is represented as $X_{i,j}$, where $i=\{1,2,...,T\}$ and $j=\{1,2,...,N\}$ denotes the patch index. Additionally, a learnable $E_{cls}$ token and a trainable position embedding $E_{pos}$ are added to each frame. With a total of $L$ transformer blocks, the input format can be represented as:
$Z^{l}=Transformer^{l}([E_{cls};{X_{i,j}^{l-1}]+E_{pos}}),$ (1)
@@
ViViT-L(Arnab et al. 2021) IN-21K 32$\times$3$\times$4 1.0$\times$12 - - 65.9 89.9
Video Swim-B(Liu et al. 2022b) IN-21K 32$\times$3$\times$1 0.3$\times$3 - - 69.6 92.7
MViT-B(Fan et al. 2021) Kinetics-600 32$\times$3$\times$1 0.2$\times$3 - - 68.7 91.5
UniFormer V2-L/14(Li et al. 2023) CLIP-400M 16$\times$3$\times$1 0.9$\times$3 60.5 86.5 72.1 93.6
ILA-ViT-L/14(Tu et al. 2023) CLIP-400M 16$\times$3$\times$4 0.9$\times$12 - - 67.8 90.5
S-ViT-B/16(Zhao et al. 2023) CLIP-400M 32$\times$3$\times$2 0.3$\times$6 - - 69.3 92.1
@@
EVL-L/14(Lin et al. 2022) CLIP-400M 8$\times$3$\times$1 0.8$\times$3 - - 65.1 -
ST-Adapter-L/14(Pan et al. 2022) CLIP-400M 16$\times$3$\times$1 1.4$\times$3 - - 71.9 93.4
DUALPATH-L/14(Park et al. 2023) CLIP-400M 16$\times$3$\times$1 0.5$\times$3 - - 70.2 92.7
AIM-L/14(Yang et al. 2023) CLIP-400M 16$\times$3$\times$1 1.9$\times$3 - - 69.4 92.3
DiST-L/14(Qing et al. 2023) CLIP-400M 16$\times$3$\times$1 1.4$\times$3 - - 72.5 93.0
M2-CLIP-B/16(Wang et al. 2024) CLIP-400M 32$\times$3$\times$1 0.8$\times$3 - - 69.1 -
Side4Video-B/16(Yao et al. 2023) CLIP-400M 16$\times$3$\times$2 0.4$\times$6 60.7 86.0 71.5 92.8
@@
Table 1: Comparison of performance with state-of-the-art methods on SSv1 & SSv2. “Cr.” and “Cl.” are the abbreviation for “spatial crops” and “temporal clips”.
Visual Temporal Encoder
SME-Adapter:
$D_{i}=Concat(V_{i-n+1}-V_{i-n},...,V_{i+n}-V_{i+n-1})$ (2)
The extracted features are processed through a 3D convolution module to align with the shape of the original features after convolution. Subsequently, feature fusion is performed via element-wise addition, resulting in the final SME-Adapter output features $S_{out}^{i}$. Formally, this stage can be represented as:
$S_{out}^{i}=\alpha Conv1(D_{i})+\beta Conv2(V_{i}),$ (3)
where, the $\alpha$ and $\beta$ represent adjustment coefficients. $S_{out}^{i}$ will be used as an input to the side network and integrated with the VIT output of the first frozen CLIP into the TDS Block.

### v2

# TDS-CLIP: Temporal Difference Side Network for Efficient Video Action Recognition

@@ Abstract
Recently, large-scale pre-trained vision-language models (e.g., CLIP), have garnered significant attention thanks to their powerful representative capabilities. This inspires researchers in transferring the knowledge from these large pre-trained models to other task-specific models, e.g., Video Action Recognition (VAR) models, via particularly leveraging side networks to enhance the efficiency of parameter-efficient fine-tuning (PEFT). However, current transferring approaches in VAR tend to directly transfer the frozen knowledge from large pre-trained models to action recognition networks with minimal cost, instead of exploiting the temporal modeling capabilities of the action recognition models themselves. Therefore, in this paper, we propose a novel memory-efficient Temporal Difference Side Network (TDS-CLIP) to balance knowledge transferring and temporal modeling, avoiding backpropagation in frozen parameter models. Specifically, we introduce a Temporal Difference Adapter (TD-Adapter), which can effectively capture local temporal differences in motion features to strengthen the model’s global temporal modeling capabilities. Furthermore, we designed a Side Motion Enhancement Adapter (SME-Adapter) to guide the proposed side network in efficiently learning the rich motion information in videos, thereby improving the side network’s ability to capture and learn motion information. Extensive experiments are conducted on three benchmark datasets, including Something-Something V1&V2, and Kinetics-400. Experimental results show that our method achieves competitive performance in video action recognition tasks. Code will be available at https://github.com/BBYL9413/TDS-CLIP.
Index Terms:
Action recognition, side network, temporal modeling, parameter-efficient fine-tuning, motion information.
I Introduction
@@
In recent years, Large Language Models (LLMs) have achieved significant success [1, 2, 3, 4, 5, 6, 7] in the fields such as text-video retrieval [8], text-to-image generation [9], video re-identification [10], etc., gradually extending their influence to the domain of computer vision. Video Action Recognition (VAR) tasks, a challenging research area within computer vision, have seen mainstream large language models represented by the CLIP paradigm [11], which includes a Vision Transformer (ViT) vision encoder [12] for image processing and a text encoder. CLIP establishes a strong association between images and text and is applied to various video tasks. This raises a thought-provoking question: how can we effectively utilize the pre-training of such large models for VAR tasks?
@@
Figure 1: Comparison of per-video GFLOPs and accuracy on SSv2 [13]. The size of the bubbles represents the memory usage during training. EVL: Efficient Video Learning [14]. ST-Adapter: Spatio-Temporal Adapter [15].
@@
Recently, the emergence of Parameter-Efficient Fine-Tuning (PEFT) methods [16, 17] has effectively alleviated this issue. PEFT adapts specific tasks by adjusting the learnability of certain parameters or introducing various adapters, avoiding extensive parameter adjustments across the entire model. With the advent of PEFT, researchers are exploring the possibility of freezing the original CLIP parameters and incorporating various adapters [18] or prompts [5]. These methods achieve impressive results in video understanding and are widely used in tasks such as video action recognition. As shown in Fig. 2(a), a popular design for efficient transfer learning is to insert tunable structures between pre-trained and parameter-frozen CLIP visual transformer blocks. Although introducing adapters under frozen parameters facilitates temporal action modeling, backpropagation still passes through the frozen layers, consuming memory. So, is there a way to solve this problem effectively?
Inspired by these works [19, 20, 21, 22], we argue that independent side-network models can better uncover temporal cues without compromising pre-training. Traditional convolutional neural networks go through a similar process of migrating pre-training from Image-Net into various variants, so we believe that CLIP can do the same. For VAR tasks, there is often a need to emphasize the effective handling of temporal branches. Previous approaches ignore the importance of independent temporal branches, and thus the potential of these approaches for temporal modeling using side networks in video understanding tasks has yet to be fully explored and exploited.
@@
Figure 2: Comparison with commonly used and effective fine-tuning methods for video recognition. (a): ST-Adapter: Spatio-Temporal Adapter [15]. (b): Ladder Side-tuning [19].
@@
In this work, we propose a method based on CLIP fine-tuning that balances memory efficiency and temporal modeling performance, called TDS-CLIP. The idea of this work is that since there are not enough computing resources to explore the entire spatial, temporal, and fusion mechanism, we rely on CLIP’s powerful spatial modeling capabilities to focus on serving temporal branches, and then combine independent temporal branches with the effective spatiotemporal modeling mechanism in traditional convolutional neural networks to form our TDS-CLIP. First, we shift the adapter positions used for the CLIP vision branch to the side network to achieve memory-efficient pre-training fine-tuning. To better model the temporal information of videos, we design a new Temporal Difference Adapter (TD-Adapter) that effectively captures local temporal differences to enhance the model’s global temporal modeling capability. Second, we integrate the TD-Adapter with a multi-head self-attention module and feed-forward network to form a learnable side network module, used for parameter training and backpropagation. Finally, to improve the side network’s capability of capturing video motion information, we introduce a Side Motion Enhancement Adapter (SME-Adapter) that acts on the input of the side network to guide it in effectively learning additional motion information relevant to action labels.
@@
We evaluate the performance of our method on several large-scale temporal action recognition datasets, including Something-Something V1 [13], Something-Something V2 [13], and Kinetics-400 [23]. Experimental results demonstrate that the proposed method effectively accomplishes video action understanding tasks.
@@
  • We propose a novel memory-efficient and temporal-branching-focused Temporal Difference Side Network for video action recognition that balances knowledge transfer and temporal modeling without requiring backpropagation in models with frozen parameters.
@@ II Related Work
VLM for Video Action Recognition. ViT [12] models are gradually replacing convolutional neural networks in the mainstream due to their powerful learning capabilities and flexibility [24, 25, 26, 27, 28, 29, 30]. EVA-CLIP [31] proposes a 64-layer ViT-E with larger scale parameters, and the effective adaptation of such a huge image model to the video domain is a task well worth exploring. Meanwhile, Transferring pre-trained models of language-images from the ViT architecture to video attracts widespread attention [32]. Previous action recognition methods rely on pre-trained models from ImageNet [33]. Currently, powerful text-image pre-trained models based on CLIP are gradually emerging and are increasingly used in VAR tasks [34]. EVL [14] uses parallel Transformer decoders to extract spatial features from the frozen CLIP model for VAR tasks. ST-Adapter [15] introduces a spatiotemporal convolution adapter to endow the CLIP model with the capability of understanding videos. AIM [35] introduces spatial adaptation, temporal adaptation, and joint adaptation, gradually equipping image models with spatiotemporal reasoning abilities, further enriching the types of adapters to suit different video action recognition tasks. GPT4Ego [36] proposes a VLM-based zero-shot egocentric action recognition method capable of exploiting the rich semantic and contextual details in self-centered videos for fine-grained conceptual and descriptive alignment, as well as designing a new selforiented visual parsing strategy thereby learning the visual contextual semantics associated with actions. DUALPATH [37] achieves dynamic modeling using spatial and temporal adaptation paths, merging successive frames into a grid-like frameset to accurately infer relationships between markers. M2-CLIP [18] introduces a TED-Adapter, enhancing the model’s generalization ability while maintaining video action recognition performance, and successfully uses it for multimodal VAR tasks. These methods effectively transfer the powerful pre-trained CLIP visual encoder to VAR tasks but do not address the issue of memory consumption during backpropagation.
@@
Side Network. The side-tuning network [38] is first proposed to address overfitting and the problem of forgetting in incremental learning, and it applies to various downstream tasks. In vision tasks, LST [19] tackles this problem by implementing a side branch network attached to the CLIP visual model. This scheme is effective in avoiding the passing of gradients in the visual encoder of frozen parameters thus reducing memory usage. Inspired by the LST, DiST [20] uses a dual encoder structure, employing the pre-trained CLIP base model as the spatial encoder and introducing a lightweight additional network as the temporal encoder, finally integrating spatiotemporal information through an ensemble branch to achieve memory-efficient video understanding. Side4Video [21] further explores the effective forms of memory-efficient side networks in video action recognition tasks and proposes a spatiotemporal side network for video understanding tasks. Although these methods demonstrate that using side networks can achieve memory-efficient image-to-video transfer learning, they do not deeply explore the temporal modeling capabilities of side network models, which we further enhance. However, we can see that these methods do not favor temporal branching and treat everything equally, thus limiting performance. We would like to liberate temporal branching completely and keep it fully expandable. Inspired by the powerful performance of convolutional neural networks [39, 40, 41] in temporal modeling, in this work, we introduce corresponding temporal modeling adapters at both the input and the path of the side network composed of ViT encoders, transforming it into a memory-efficient and powerful temporal modeling framework for VAR tasks. Its extensibility in temporal modeling is well preserved and its ability to reduce memory overhead by exploiting the characteristics of the measurement network provides some meaningful assistance to subsequent researchers.
III Method
@@
As shown in Fig. 3, the overall architecture of TDS-CLIP framework consists of two key components: the visual spatial encoder and the visual temporal encoder. The visual spatial encoder is composed of a frozen CLIP pre-trained ViT, which provides strong spatial semantic features for the input video frames. The visual temporal encoder is a side network structure that accepts the output features from the visual spatial encoder. It takes the dense video frame features provided by the SME-Adapter as input and transforms the spatial features from the frozen pre-trained CLIP branch into an effective form suitable for temporal modeling in video understanding.
III-A Denotations
@@
The spatial encoder consists of the ViT with frozen parameters from the CLIP visual branch, which can independently capture the spatial semantic features of sparse input video frames, thereby leveraging its powerful pre-training to enhance training performance. Following the ViT format, the input frames are first divided into $N=\frac{HW}{P^{2}}$ patches, each with a shape of $P^{2}$, and then projected through a fully connected layer. Then, for the input full-length video frame sequence, each patch feature vector is represented as $X_{i,j}$, where $i=\{1,2,...,T\}$ and $j=\{1,2,...,N\}$ denotes the patch index. Additionally, a learnable $E_{cls}$ token and a trainable position embedding $E_{pos}$ are added to each frame. Mathematically, the frame-level input is constructed as:
$[e^{(0)},x^{(0)}]=([E_{cls};{X_{i,j}]+E_{pos}}),$ (1)
with a total of $L$ transformer blocks, if we place the visual encoder before each transformer layer, the inputs will be processed sequentially, this format can be represented as:
$Z^{l}=Transformer^{l-1}([e^{(l-1)},x^{(l-1)}]),$ (2)
III-B Visual Temporal Encoder
@@
III-B1 SME-Adapter
$D_{i}=Concat(V_{i-n+1}-V_{i-n},...,V_{i+n}-V_{i+n-1}),$ (3)
subsequently, the SME-Adapter output feature $S_{out}^{i}$ is fused into the original ViT structure in the form of residuals, while introducing factor control weights. $S_{out}^{i}$ will be used as an input to the side network and integrated with the ViT output of the first frozen CLIP into the TDS block. Formally, this phase can be represented as:
$S_{out}^{i}=\alpha Reshape(Conv(D_{i}))+\beta\mathrm{ViT}(V_{i}),$ (4)
where, the $\alpha$ and $\beta$ represent adjustment coefficients and the output after convolution needs to be reshaped into the form of ViT.
III-B2 TD-Adapter
As shown in Fig. 3(b), the input feature of the TD-Adapter is denoted as $Z_{v}\in\mathbb{R}^{T\times N\times C}$. First, it is reshaped to a form $Z_{s}\in\mathbb{R}^{C\times T\times H\times W}$ suitable for 3D convolution. Then, a set of convolution operations is used for dimensionality reduction to obtain the feature $Z_{s}\in\mathbb{R}^{C/r\times T\times H\times W}$, $r$ denotes rate of reduction. Next, we capture the sensitive feature values of the local temporal dimension features using a 3D max-pooling operation, and a feature difference operation is employed to obtain dynamic features of local motion patterns. Compared to conventional displacement [39], pooling is a simpler and more effective operation.
$D_{z}=Z_{s}-Pool(Z_{s}),$ (5)
where $D_{z}$ denotes the feature difference. This method is commonly used in the early stages of convolutional models for efficient temporal modeling [42, 39, 40].
III-C TDS-CLIP for video action recognition
We integrate the proposed SME-Adapter and TD-Adapter into Side4Video [21] architecture, where the TD-Adapter, combined with a trainable ViT encoder, forms the TDS-CLIP block. We retain the CLS token shift  [43, 44] operation to further improve the capacity for learning temporal information, offering the advantage of minimal additional memory overhead. The number of these blocks matches the frozen ViT encoders in the visual branch of CLIP and combined via side-connections. For action recognition tasks, we obtain the prediction scores by performing global average pooling on the output of the final TDS-CLIP block.
TABLE I: Comparison of performance with state-of-the-art methods on SSv1 & SSv2. “Cr.” and “Cl.” are the abbreviation for “spatial crops” and “temporal clips”.
@@ III-C TDS-CLIP for video action recognition
ViViT-L [45] IN-21K 32$\times$3$\times$4 1.0$\times$12 - - 65.9 89.9
Video Swim-B [46] IN-21K 32$\times$3$\times$1 0.3$\times$3 - - 69.6 92.7
MViT-B [47] Kinetics-600 32$\times$3$\times$1 0.2$\times$3 - - 68.7 91.5
UniFormer V2-L/14 [48] CLIP-400M 16$\times$3$\times$1 0.9$\times$3 60.5 86.5 72.1 93.6
ILA-ViT-L/14 [49] CLIP-400M 16$\times$3$\times$4 0.9$\times$12 - - 67.8 90.5
S-ViT-B/16 [50] CLIP-400M 32$\times$3$\times$2 0.3$\times$6 - - 69.3 92.1
STAN-B/16 [51] CLIP-400M 16$\times$3$\times$1 0.7$\times$1 - - 69.5 92.7
@@
EVL-L/14 [14] CLIP-400M 8$\times$3$\times$1 0.8$\times$3 - - 65.1 -
ST-Adapter-L/14 [15] CLIP-400M 16$\times$3$\times$1 1.4$\times$3 - - 71.9 93.4
DUALPATH-L/14 [37] CLIP-400M 16$\times$3$\times$1 0.5$\times$3 - - 70.2 92.7
AIM-L/14 [35] CLIP-400M 16$\times$3$\times$1 1.9$\times$3 - - 69.4 92.3
DiST-L/14 [20] CLIP-400M 16$\times$3$\times$1 1.4$\times$3 - - 72.5 93.0
M2-CLIP-B/16 [18] CLIP-400M 32$\times$3$\times$1 0.8$\times$3 - - 69.1 -
Side4Video-B/16 [21] CLIP-400M 16$\times$3$\times$2 0.4$\times$6 60.7 86.0 71.5 92.8
@@
TDS-CLIP-L/14 CLIP-400M 16$\times$3$\times$2 1.6$\times$6 64.1 88.2 74.1 94.0
III-D Training Loss
Due to the removal of the text branch, the visual-only task requires that the model can focus more on spatio-temporal modeling of the video action, and thus can be optimized directly using the cross-entropy loss function. Given an input video, the side network generates a video representation $f_{out}\in\mathbb{R}^{T\times N\times d}$, where $d$ denotes the final linear layer mapping output shape, and then applies Global Average Pooling (GAP) to the patch markers to obtain the final representation $f_{emb}$. This process is denoted as:
$f_{emb}=GAP(Proj(f_{out})),$ (6)
in order to mitigate model overfitting and improve generalization across classes, we introduce label-smooth to regularize constraints on the loss function. The final shape of the loss function combining the label smoothing method is:
$\mathcal{L}_{\text{CE-LS}}=\left[(1-\alpha)\cdot\mathbb{I}(i=y)+\alpha/N\right]\mathcal{L}_{\text{CE}},$ (7)
where $\alpha$ is the smoothing intensity, $\mathbb{I}()$ is the exponential function, $y$ is the video category label, and $N$ is the total number of categories used for classification in the dataset. Further, the final loss function representation for training can be obtained by expanding $\mathcal{L}_{\text{CE}}$, where $Y_{i}$ denotes the category with label smoothing regularization.
$\mathcal{L}_{\text{CE}}=-\sum_{i=1}^{N}Y_{i}\log\left(\frac{exp(f_{emb},i)}{\sum_{j=1}^{N}exp(f_{emb},j)}\right)$ (8)
IV Experiments
IV-A Experimental Setup
To evaluate the effectiveness of our method, we assess it on several widely used action recognition benchmark datasets, including Something-Something V1 (SSv1) [13], Something-Something V2 (SSv2) [13], and Kinetics400 (K400) [23]. SSv1 and SSv2 are commonly used datasets for evaluating the importance of temporal modeling. Kinetics400 is a comprehensive action recognition dataset covering 400 different human actions and is used to evaluate the generalization performance of the methods. In Kinetics400, most action categories are biased toward static scene context. In SSv1 and SSv2, the action categories are less related to the static scene context, but are closely related to the dynamic information in the video.
As in the previous method [15, 18], we use the pre-trained CLIP ViT-B/16 and ViT-L/14 as spatial encoders. By adjusting the dimensions, our model balances memory usage and performance. Due to hardware dilemmas, unless otherwise specified, we use the corresponding ViT-B/16 or ViT-L/14 structure in the temporal encoder to form the TDS-CLIP framework, employing a sparse frame sampling strategy of 8 or 16 frames during both training and inference. For more details on parameter settings and techniques, please see the Supplementary Material.
TABLE II: Comparison of performance with state-of-the-art methods on Kinetics-400. “Cr.” and “Cl.” are the abbreviation for “spatial crops” and “temporal clips”.